OpenStudy (anonymous):

Quick algebra help 2x+-40=-6 I really dont understand this problem, we're suppose to find what X is, however, this cant be true! -40 - -6=-46, X=-23? how is this true! I mean, the only wrong thing I could of done is if X IS a negative number, in which case when divided, gives a positive number...but still..

OpenStudy (anonymous):

So you can rewrite htis as: 2x - 40 = -6, now we need to isolate the x so we can add 40 to each side, this gives us 2x = 40 -6 = 34, now to isolate x we divide across by 2 and get x = 17.

OpenStudy (anonymous):

Waaaaaaooo, I see what you did there, thanks man.